OpenStudy (anonymous):

pls. help me solve this. find the derivatives y=x/x^2-1 f(x)=x^2+1/x^2-1 pls show your solution.

OpenStudy (akshay_budhkar):

the working is..

OpenStudy (anonymous):

how is it done?

OpenStudy (akshay_budhkar):

use the division rule of derivatives to get \[{(x^2-1)(1)-(x)(2x)}/(x^2-1)^2\]

OpenStudy (akshay_budhkar):

for the first one.. the answer is -(x^2+1)/(x^2-1)^2

OpenStudy (akshay_budhkar):

do you get it? or do u want me to explain more?

OpenStudy (akshay_budhkar):

shessa???? do you know the formula?

OpenStudy (anonymous):

\[v*du/dx- u*dv/dx/v^2\]

OpenStudy (anonymous):

pls explain further.

OpenStudy (akshay_budhkar):

right.. so i have applied the same formula to get the answer..:)

OpenStudy (akshay_budhkar):

then you solve the bracket to get your answer.. you want me to do it too?

OpenStudy (anonymous):

yup pls.

OpenStudy (akshay_budhkar):

(v∗du/dx−u∗dv/dx)/v2 its this the whole under v^2..

OpenStudy (akshay_budhkar):

oh ok you get in the numerator x^2-1-2x^2...

OpenStudy (anonymous):

pls go to this website http://www.twiddla.com/584540

OpenStudy (anonymous):

x^2-1(1)=X^2-1 then x(2x)=?

OpenStudy (akshay_budhkar):

sorry shessha had a network problem.. you there????

OpenStudy (anonymous):

yup i'm still here

OpenStudy (anonymous):

ashay pls go to this website http://www.twiddla.com/584540

OpenStudy (akshay_budhkar):

yes come on twiddla